John can paint the house in 4 hours. Mike can paint the same house in 6 hours. How long would it take both John and mike to paint the same house together.
Answer by stanbon(48558)

John can paint the house in 4 hours. Mike can paint the same house in 6 hours. How long would it take both John and mike to paint the same house together.
-----------------
John DATA:
Time = 4 hrs/job ; Rate = 1/4 job/hr.
-----------------------
Mike DATA:
Time = 6 hrs/job ; Rate = 1/6 job/hr.
-----------------------
Together DATA:
Time = x hrs/job ; Rate = 1/x job/hr.
-----------------------
EQUATION:
rate + rate = together rate
1/4 + 1/6 = 1/x
Multiply thru by 12x to get:
3x + 2x = 12
5x = 12
x = 12/5 = 2 hrs 24 minutes
=========================
Cheers,
Stan H.